Solve for x :
2x/3 + 1/4 = x/3 - 1/12

Answer by Cromlix(4381)

2x/3 + 1/4 = x/3 - 1/12
Use common denominator = 12
8x + 3 = 4x - 1
8x - 4x = -1 -3
4x = - 4
x = -1
Check answer using 2x/3 + 1/4 = x/3 - 1/12
with x = -1
2(-1)/3 + 1/4 = -1/3 - 1/12
-2/3 + 1/4 = -1/3 - 1/12
-5/12 = -5/12 Proved